OpenStudy (anonymous):

The water faucet can fill a tub in 15 minutes. But there is a whole in the tub through which a full tub of water can drain out in 18 minutes. If the tub is empty and the faucet is turned on, how long will it take to fill the tub?

OpenStudy (anonymous):

@mary.rojas

OpenStudy (anonymous):

talk to me ineed your help soory

OpenStudy (anonymous):

plz i need your help

OpenStudy (anonymous):

post your question

OpenStudy (anonymous):

in 1 min tub is filled by faucet =1/15 in 1min water drained out from tub by hole=1/18 if both are working at the same time then water filled in the tub in 1 min=(1/15)-(1/18)=?

OpenStudy (anonymous):

oh this is a work problem thanks!!

OpenStudy (jdoe0001):

the faucet is filling it up in 15mins the hole is draining it in 18mins so that means the faucet is going at a faster rate, since it takes less time to fill it up

OpenStudy (anonymous):

wait so would it be (1)/(1/15-1/18)=x ?

OpenStudy (jdoe0001):

so if you check how much each does in 1 minute, you'd get the rate at which is filling up

OpenStudy (anonymous):

so am i right ?

OpenStudy (anonymous):